OpenStudy (anonymous):

Consider the solid obtained by rotating the region bounded by the given curves about the x-axis. y = 5 - (5/2)x y = 0 x = 0 , x = 1 find the volume

OpenStudy (dumbcow):

\[V = \pi \int\limits_{0}^{1} (5-\frac{5}{2}x)^{2} dx\] \[=\pi \int\limits_{0}^{1}\frac{25}{4}x ^{2}-25x+25 dx\] \[=\pi[\frac{25}{12}x ^{3} -\frac{25}{2}x ^{2}+25x] \left\{ 0 \to 1 \right\}\] \[=\frac{175}{12}\pi\]

OpenStudy (anonymous):

omg thaats what i got, i just forgot to add the pi sign lol. thank you so much :)

OpenStudy (dumbcow):

haha your welcome